# Retrieval of photon statistics from click detection

### v1

@@ Abstract
We utilize click-counting theory for the reconstruction of photon statistics. Our approach employs an analytic pseudo-inversion method to estimate photon counts from measured click counts. A reconfigurable time-bin multiplexing, click-counting detector is set up that renders it possible to alter the photon-number resolution as needed. A detector tomography is carried out, yielding vital measurement features, such as quantum efficiencies, cross-talk rates, etc. We gauge the success of the pseudo-inversion by applying the Mandel and binomial parameters, resulting in an additional interpretation of these parameters for the discrimination of distinct quantum statistics. In addition, we apply a loss deconvolution technique to account for detection losses.
@@ I Introduction
Since the beginning of quantum mechanics, the field of quantum optics has attracted an extraordinary interest for purposes of explaining optical phenomena that are outside the realm of classical physics [1, 2, 3]. Since then, ongoing research has employed nonclassical optical states [4, 3] to develop and advance photonic quantum technologies [5, 6, 7, 8]. Examples of applications are boson sampling [9], quantum key distribution [10], the study of quantum interference [11, 12], quantum metrology [13, 14], and coherent control for quantum-state engineering [15, 16, 17], all of which pertain to the fields of quantum information [18, 19] and communication [20, 21].
@@
For understanding and manipulating quantum light, a prerequisite is the ability to measure quantum-optical systems [22]. The information obtained from these measurements allows us to quantify the resourcefulness of quantum states by reconstructing their full interference structure [23, 24, 25]. One of the most crucial questions that come up in this context is determining whether a given light field is quantum or not [26, 27, 28, 29], often investigated via photon statistics [30, 31], and photo-detectors are the main tool to provide this information and thus are vital elements in quantum state reconstruction.
@@
However, true photon-number-resolving detectors, which can measure arbitrarily high photon counts, are relatively inaccessible, despite ample research efforts. By contrast, detectors that record clicks as a result of any number of photons are readily available for on-demand usage due to their technological abundance [32, 33, 34, 35, 36, 37]. Therefore, it is of utmost importance to tailor strategies that account for the nature of click detection when characterizing nonclassical photon statistics.
@@
The quantum-optical click-detection theory allows one to formulate a proper model of click-counting systems, which may include up to one hundred individual click detectors [38, 39, 40, 41, 42, 43]. To this end, experimental methods, such as multiplexing [44, 45, 46, 47, 48], are often used. An operator version of a binomial probability distribution arises as the general description of click-counting devices [49]. Thus, the earlier notion of sub-Poissonian light for photon-number statistics [50] is updated to the concept of sub-binomial light as its click-based equivalent [51], being further supplemented by the concepts of sub-Poisson-binomial light and sub-multinomial light [52, 53]. For identifying sub-binomial light, the binomial parameter $Q_{B}$ can be experimentally applied [54, 55], just like the Mandel parameter $Q_{M}$ is for photon-number statistics.
@@
The active and vast research in the broad area of reconstruction includes a manifold of sophisticated techniques to obtain photon-number distributions from measured data recorded with contemporary detection systems [56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. These reconstruction tools include maximum likelihood estimation [66], detector tomography [67], and regularizations of ill-posed problems [68], to name but a few. Another noteworthy method within the context of this work is a detector calibration to determine the detector’s response to light (see, e.g., Refs. [69, 70]), which allows one to determine important measurement characteristics, e.g., light-matter interactions [71], noise [72], and dead time and after pulses [73]. Regardless, the aforementioned methods commonly require a detection scheme with high efficiency, low noise, signals far from saturation, etc., being hardly satisfiable in practice.
@@
In this work, we extract the photon statistics from measured click-counting statistics by utilizing an easy-to-implement pseudo-inversion technique, theoretically proposed in Ref. [74], circumventing the complexities of other reconstruction techniques. Furthermore, the Mandel and binomial parameters are employed unconventionally as discriminators between click-type and photon-number-type quantum statistics, thereby assessing the success of the pseudo-inversion procedure for calibration measurements with coherent light. A time-bin multiplexing detection system with superconducting single-photon nanowire detectors is realized in a modular manner, allowing us to alter the photon-number-resolution of the click-detection system as needed. A finely-grained detector calibration is carried out to assess vital parameters of the detection system. Eventually, we also apply a loss deconvolution method to retrieve a near-ideal photon-number distribution from our data.
@@
The reminder of the paper is organized as follows: The distinct photon-number and click-counting theories and their discrimination via $Q$ parameters are laid out in Sec. II. In Sec. III, we describe our experiment and present the reconstruction of the detector response function via coherent-state measurements. Section IV includes our key results from the pseudo-inversion and the application of $Q$ parameters. The losses are treated in Sec. V. Eventually, we conclude in Sec. VI.
II Detection theories
This section covers the theoretical fundamentals of click-counting and photon-number statistics and measurements. We elaborate on how one can study quantum light when having access to click counts only. Also, parameters are reviewed for determining the Poissonian and binomial character of probability distributions.
II.1 Photon and click statistics
In photoelectric detection theory [75, 76], ideal photodetectors can be represented through their positive operator-valued measure,
$\hat{\Pi}_{n}=|n\rangle\langle n|={:}e^{-\hat{n}}\frac{\hat{n}^{n}}{n!}{:},$ (1)
@@ II.1 Photon and click statistics
$\hat{\Gamma}=\eta\hat{n}+\nu\hat{1},$ (2)
where $\eta$ is the quantum efficiency and $\nu$ describes dark counts. Photon counting, as described through Eq. (1) together with general response functions $\hat{\Gamma}$, always takes the form of quantum-Poissonian distribution, i.e., an operator-based Poissonian formula. To describe the quantum-statistical properties of states of light, access to the photon-number distribution is essential, which is difficult because of a lack of true photon-number-resolving detectors. Thus, it is paramount to explore techniques to infer photon-number information from data measured with state of the art detection systems.
Figure 1: The click-counting theory applies to, for example, the three depicted multiplexing schemes, in which light is uniformly split spatially ($50{:}50$ beam splitters) and temporally (delay loops), before being measured with on-off click detectors that collectively yield a total number of clicks. The top, spatial multiplexing layout utilizes three beam splitters and $N=4$ detectors. The middle configuration uses one less beam splitter and only two detectors. Together with the delay line, this results in early and late time bins in the top and bottom detector, i.e., $N=4$ detection bins. The bottom scheme doubles the number of time bins by introducing only one extra beam splitter and a delay double the length of the prior loop, resulting in a $N=8$ time-bin multiplexing detector. The latter two schemes, being more efficient than the spatial multiplexing in the number of optical elements, are experimentally realized in this work.
Nowadays, common detector technologies employ so-called click detectors where a “click” is recorded if one or an arbitrary number of photons are absorbed, and “no click”’ is recorded otherwise [32, 34, 36, 37]. Thereby, one defines ideal measurement operators as follows:
$\hat{\pi}_{0}=|0\rangle\langle 0|={:}e^{-\hat{n}}{:},\quad\hat{\pi}_{1}=\hat{1}-\hat{\pi}_{0}=\sum_{n=1}^{\infty}|n\rangle\langle n|,$ (3)
where the subscript 00 and $1$ denote no click and one click, respectively. As for the photon-counting detectors, we can replace $\hat{n}$ with $\hat{\Gamma}$ to model detector imperfections. Furthermore, incident light may be uniformly distributed across $N$ click detectors, commonly referred to as multiplexing [44, 47, 45, 46]; see Fig. 1. Then, we can express realistic click detectors through [49]
@@
for $k\in\{0,\ldots,N\}$ total clicks. One strategy of multiplexing is time-bin multiplexing, which is the detection scenario implemented in our experiment (two bottom schemes in Fig. 1). In contrast to the photon-number distribution, the click-counting distribution [Eq. (4)] is a quantum-operator analog to a binomial distribution. For example, the lossless and dark-count-free click-counting distribution for coherent light is $c_{k}=\langle\alpha|\hat{\pi}_{k}|\alpha\rangle=\binom{N}{k}\big{(}e^{-|\alpha|^{2}/N}\big{)}^{N-k}\big{(}1-e^{-|\alpha|^{2}/N}\big{)}^{k}$. Also, the Poissonian distribution is approached in the limit of $N\to\infty$ detection bins [49].
It is important to emphasize that $k$ clicks and $n$ photons are related, but dissimilar concepts (even if $k=n$, $\eta=1$, and $\nu=0$), which is also discussed later in more detail. Furthermore, since it is not assumed that each click comes from one photon, Eq. (4) applies to arbitrarily high intensities, thus naturally accounting for detector saturation.
II.2 Quantum-statistical signatures
One can use the Poissonian behavior in the photoelectric detection scenario as a reference for probing nonclassicality of light [50]—if one has access to the photon-number. Analogously, one can characterize the super- and sub-binomial behavior with click detection systems for determining the quantumness of the signal in theory and experiment [51, 54]. In photoelectric measurements, the Mandel parameter is utilized,
@@ II.2 Quantum-statistical signatures
Introducing the parameter $Q_{B}$ is necessary because if one applies $Q_{M}$ to click-based measurement devices, we may get a false indication of nonclassicality [49, 51]. Here, we additionally attribute another function to $Q_{B}$ and $Q_{M}$ as they enable us to determine the Poissonian or binomial behavior of a detection system and the character of reconstructed distributions from data; see Fig. 2. In particular, we are going to reconstruct parts of the photon-number distribution from click-counting data for coherent light by employing a pseudo-inversion [74] and probe the success of this procedure by answering the question: $Q_{M}=0$?
@@ III Setup and detector calibration
In this section, we describe and characterize our experiment. For instance, the detector characteristics of the whole measurement system are analyzed through a thorough detector tomography. Beyond previous characterizations, we also probe the cross-talk and splitting uniformity between the detection time-bins.
@@ III.1 Setup description
We begin with a description of the realized detection system; see Fig. 3. We use a pulsed laser with a pulse length of $100\,\mathrm{ps}$ at $\lambda=1549.8\,\mathrm{nm}$ wavelength as a coherent light source. The repetition rate of the pulsed laser is $r=2\,\mathrm{MHz}$. The pulse intensity is split via a beam splitter. The bright output, containing approximately $94\%$ of the light, is directly measured with a power meter and is used as a reference to monitor power drifts. The remaining $6\%$ is attenuated to the single-photon level with two variable optical attenuators. The two attenuators have a rated error of ±$\pm 0.3\,\mathrm{dB}$ and are individually characterized for linearity. The attenuated coherent laser pulses are connected to the time-bin multiplexing scheme, which can be set to consist of either $N=4$ or $N=8$ detection bins. The output from the time multiplexing is detected with two superconducting nanowire single-photon detectors. A time-to-digital converter is used to log the arrival time of the detected clicks relative to the laser pulses.
Before the measurements are performed, the input photon flux to the detectors is calibrated. For that, the power exiting the two output ports of the variable beam splitter are measured with the power meter. Additionally, the zero attenuation of the attenuators is measured and is added as a constant offset to the selected attenuation. Together with the constant power monitoring of the input power, this calibration provides the photon flux entering the time-bin multiplexing detection system within the measurement uncertainty.
@@ III.2 Characterization
Figure 4: Top plot depicts the mean clicks per bin, Eq. (9), depicted here for the $N=4$ configuration and the maximal photon number $\bar{n}=|\alpha|^{2}=0.84$. This distribution is compatible with a uniform distribution (purple, solid line); one standard deviation $\sigma$ of ±$\pm 0.011$ clicks (blue, dashed lines) is shown. Bottom plot depicts bin-bin cross-talk through vanishing off-diagonal covariances, Eq. (10), for the same data set as used above. No two bins $j\neq j^{\prime}$ exhibit nonzero correlations within the margin of error (blue cylinder depicts $10\sigma$ for improved visibility). Similar results are found for all measured cases, including the $N=8$-bin scenario.
Bin uniformity and cross-talk.
We can now analyze the detection system. We begin with observing the system’s quality across individual detection bins. To this end, clicks $c_{k_{1},\ldots,k_{N}}$ were recorded with $k_{j}=0$ and $k_{j}=1$ denoting no click or a click from the $j$th time bin. One measure for how well our system operates is the uniform distribution of light across detection bins, which can be quantitatively assessed with
@@
where the scaling $\hat{n}/N$ accounts for the fact that input photons are distributed across the $N$ detection bins and the second-order term proportional to $\gamma$ probes the presence of nonlinear absorption properties [77, 48]. Together with Eq. (11), we thus have the function
@@
Figure 5: Determination of the response function [Eq. (13); blue, dashed lines] from measured mean click numbers [Eq. (11); thick, purple points] for $N=4$ and $N=8$ time-bin detection setups in the top and bottom depiction, respectively. In both cases, the response is governed by a linear behavior, with quantum efficiencies $\eta|_{N=4}=60.8\%$ and $\eta|_{N=8}=60.5\%$. Dark counts (less than $2\times 10^{-4}$) and nonlinear contributions ($\gamma\bar{n}^{2}/N^{2}<10^{-3}$) are negligible in both cases. For $N=4$ and $N=8$, the coefficient of determination is $R^{2}>0.9999$, quantitatively showing a near-perfect agreement between data and fit.
We use a nonlinear regression in Fig. 5 to retrieve the coefficients in Eq. (13). For both detector variants, $N=4$ and $N=8$, the dark count contributions are negligible, $\nu|_{N=4}=1.8\times 10^{-4}$ and $\nu|_{N=8}=5.8\times 10^{-6}$. Also, the nonlinear contribution that scales with $(\bar{n}/N)^{2}$ is small, $\gamma|_{N=4}=1.9\times 10^{-2}$ and $\gamma|_{N=8}=6.2\times 10^{-2}$. Most importantly, the quantum efficiency of the click-counting device is about $60\%$, $\eta|_{N=4}=60.8\%$ and $\eta|_{N=8}=60.5\%$, including both optical losses in the multiplexing stage and non-unit efficiencies of individual on-off detectors.

### v2

# Experimental retrieval of photon statistics from click detection

@@ Abstract
We utilize click-counting theory for the reconstruction of photon statistics, employing an analytic pseudo-inversion method. A reconfigurable time-bin multiplexing, click-counting detector is experimentally implemented, allowing for altering the photon-number resolution. A detector tomography is carried out, yielding vital measurement features, such as quantum efficiencies and cross-talk rates. We gauge the success of the pseudo-inversion by applying the Mandel and binomial parameters, resulting in an additional interpretation of these parameters for the discrimination of distinct quantum statistics. For coherent states, which lie at the classical-nonclassical boundary, both parameters are highly sensitive measures, probing the kind of statistics and the reconstruction performance. In addition, we apply a loss deconvolution technique to account for detection losses.
@@ I Introduction
Since the beginning of quantum mechanics, the field of quantum optics has attracted an extraordinary interest for purposes of explaining optical phenomena that are outside the realm of classical physics [1, 2, 3]. Since then, ongoing research has employed nonclassical optical states [4, 3] to develop and advance photonic quantum technologies [5, 6, 7]. For understanding and manipulating quantum light, a prerequisite is the ability to measure quantum-optical systems [8]. The information obtained from these measurements allows us to quantify the resourcefulness of quantum states by reconstructing their full interference structure [9, 10, 11]. One of the most crucial questions is determining whether a given light field is quantum [12, 13, 14], often investigated via photon statistics [15, 16] for which photo-detectors are the main tool to provide this information.
@@
However, true photon-number-resolving detectors, which can measure arbitrarily high photon counts, are relatively inaccessible, despite ample research efforts. By contrast, detectors that record clicks as a result of any number of photons are readily available for on-demand usage due to their technological abundance [17, 18, 19, 20, 21]. Therefore, it is of utmost importance to tailor strategies that account for the nature of click detection when characterizing nonclassical photon statistics.
@@
The quantum-optical click-detection theory allows one to formulate a proper model of click-counting systems, which may include up to 100 individual click detectors [22, 23, 24, 25, 26]. To this end, experimental methods, such as multiplexing [27, 28, 29, 30], are often used. An operator version of a binomial probability distribution arises as the general description of click-counting devices [31]. Thus, the earlier notion of sub-Poissonian light for photon-number statistics [32] is updated to the concept of sub-binomial light as its click-based equivalent [33], being further supplemented by the concepts of sub-Poisson-binomial light and sub-multinomial light [34, 35]. For identifying sub-binomial light, the binomial parameter $Q_{B}$ can be experimentally applied [36, 37], just like the Mandel parameter $Q_{M}$ is for photon-number statistics.
@@
A manifold of sophisticated techniques to obtain photon-number distributions from measured data has been developed [38, 39, 40, 41, 42, 43, 44, 45, 46]. These reconstruction tools include maximum likelihood estimation [47], detector tomography [48], and regularizations of ill-posed problems [49], to name but a few. Another noteworthy method within the context of this paper is a detector calibration to determine the detector’s response to light (see, e.g., Refs. [50, 51]), yielding important measurement characteristics, e.g., light-matter interactions [52], noise [53], and dead time and after pulses [54]. Regardless, the aforementioned methods commonly require a detection scheme with high efficiency, low noise, signals far from saturation, etc., being hardly satisfiable in practice.
@@
In this paper, we extract the photon statistics from measured click-counting statistics by utilizing an easy-to-implement pseudo-inversion technique [55], circumventing the complexities of other techniques. Furthermore, the Mandel and binomial parameters are employed unconventionally as discriminators between click-type and photon-number-type quantum statistics, assessing the success of the pseudo-inversion procedure. By applying coherent light, which lies at the challenging classical-quantum boundary, our procedure removes a possible loophole which could lead to fake nonclassicality. A time-bin multiplexing detection system with superconducting single-photon nanowire detectors is realized in a modular manner, and a finely-grained detector calibration is carried out. Eventually, we also apply a loss deconvolution method to retrieve a near-ideal photon-number distribution.
@@
The remainder of the paper is organized as follows: The distinct photon-number and click-counting theories and their discrimination via $Q$ parameters are laid out in Sec. II. In Sec. III, we describe our experiment and present the reconstruction of the detector response function via coherent-state measurements. Section IV includes our key results from the pseudo-inversion and the application of $Q$ parameters. The losses are treated in Sec. V. Eventually, we conclude in Sec. VI.
II Detection theories and quantum-statistical signatures
@@
Figure 1: The click-counting theory applies to, for example, the three depicted multiplexing schemes, in which light is uniformly split spatially ($50{:}50$ beam splitters) and temporally (delay loops), before being measured with on-off click detectors that collectively yield a total number of clicks. The top, spatial multiplexing layout utilizes three beam splitters and $N=4$ detectors. The middle configuration uses one less beam splitter and only two detectors. Together with the delay line, this results in early and late time bins in the top and bottom detector, i.e., $N=4$ detection bins. The bottom scheme doubles the number of time bins by introducing only one extra beam splitter and a delay double the length of the prior loop, resulting in a $N=8$ time-bin multiplexing detector. The latter two schemes, being more efficient than the spatial multiplexing in the number of optical elements, are experimentally realized in this paper.
Nowadays, common detector technologies employ so-called click detectors where a “click” is recorded if one or an arbitrary number of photons are absorbed, and “no click”’ is recorded otherwise [17, 18, 20, 21]. Furthermore, incident light may be uniformly distributed across $N$ click detectors, commonly referred to as multiplexing [27, 29, 28]; see Fig. 1. Then, we can express realistic click detectors through [31]
@@
for $k\in\{0,\ldots,N\}$ total clicks. One strategy of multiplexing is time-bin multiplexing, which is the detection scenario implemented in our experiment (two bottom schemes in Fig. 1). In contrast to the photon-number distribution, the click-counting distribution [Eq. (1)] is a quantum-operator analog to a binomial distribution. For example, the lossless and dark-count-free click-counting distribution for coherent light is $c_{k}=\langle\alpha|\hat{\pi}_{k}|\alpha\rangle=\binom{N}{k}\big{(}e^{-|\alpha|^{2}/N}\big{)}^{N-k}\big{(}1-e^{-|\alpha|^{2}/N}\big{)}^{k}$.
It is important to emphasize that $k$ clicks and $n$ photons are related, but dissimilar concepts (even if $k=n$, $\eta=1$, and $\nu=0$), which is discussed later in more detail.
One can use the Poissonian behavior in the photoelectric detection scenario as a reference for probing nonclassicality of light [32]. Analogously, one can characterize the super- and sub-binomial behavior with click detection systems for determining the quantumness of the signal in theory and experiment [33, 36]. In photoelectric measurements, the Mandel parameter is utilized,
@@
Introducing the parameter $Q_{B}$ is necessary because if one applies $Q_{M}$ to click-based measurement devices, we may get a false indication of nonclassicality [31, 33].
Here, we additionally attribute another function to $Q_{B}$ and $Q_{M}$ as they enable us to determine the Poissonian or binomial behavior of a detection system and the character of reconstructed distributions from data, Fig. 2. In general, $Q_{B/M}<0$ and $Q_{B/M}>0$ means that the distribution under consideration is, respectively, too narrow and too broad for a binomial or Poissonian distribution. Also, coherent states, as employed in our experiment, are at the classical-quantum boundary, exactly yielding $Q_{B}=0$ and $Q_{M}=0$. Thus, such edge states present a highly sensitive testbed for our method, for example, to ensure that no fake quantum effects are introduced via the reconstruction procedure.
@@ III Setup and detector calibration
In this section, we describe and characterize our experiment. The detector characteristics of the whole measurement system are analyzed through a thorough detector tomography. We also probe the cross-talk and splitting uniformity between the detection time-bins.
@@ III.1 Setup description
We begin with a description of the realized detection system; see Fig. 3. We use a pulsed laser with a pulse length of $100\,\mathrm{ps}$ at $\lambda=1549.8\,\mathrm{nm}$ wavelength as a coherent light source. The repetition rate of the pulsed laser is $r=2\,\mathrm{MHz}$. The pulse intensity is split via a beam splitter. The bright output, containing approximately $94\%$ of the light, is directly measured with a power meter and is used as a reference to monitor power drifts. The remaining $6\%$ is attenuated to the single-photon level with two variable optical attenuators. The two attenuators have a rated error of ±$\pm 0.3\,\mathrm{dB}$ and are individually characterized for linearity. The attenuated coherent laser pulses are connected to the time-bin multiplexing scheme, which can be set to consist of either $N=4$ or $8$ detection bins. The output from the time multiplexing is detected with two superconducting nanowire single-photon detectors. A time-to-digital converter is used to log the arrival time of the detected clicks relative to the laser pulses.
Before the measurements are performed, the input photon flux to the detectors is calibrated. For that, the power exiting the two output ports of the variable beam splitter is measured with the power meter. Additionally, the zero attenuation of the attenuators is measured and is added as a constant offset to the selected attenuation. Together with the constant power monitoring of the input power, this calibration provides the photon flux entering the time-bin multiplexing detection system within the measurement uncertainty.
@@ III.2 Characterization
Figure 4: The top plot depicts the mean clicks per bin, Eq. (6), depicted here for the $N=4$ configuration and the maximal photon number $\bar{n}=|\alpha|^{2}=0.84$. This distribution is compatible with a uniform distribution (purple, solid line); one standard deviation $\sigma$ of ±$\pm 0.011$ clicks (blue, dashed lines) is shown. The bottom plot depicts bin-bin cross talk through vanishing off-diagonal covariances, Eq. (7), for the same data set as used above. No two bins $j\neq j^{\prime}$ exhibit nonzero correlations within the margin of error (blue cylinder depicts $10\sigma$ for improved visibility). Similar results are found for all measured cases, including the $N=8$-bin scenario.
Bin uniformity and cross talk.
We can now analyze the detection system. We begin with observing the system’s quality across individual detection bins. To this end, clicks $c_{k_{1},\ldots,k_{N}}$ were recorded with $k_{j}=0$ and $1$ denoting no click or a click from the $j$th time bin. One measure for how well our system operates is the uniform distribution of light across detection bins, which can be quantitatively assessed with
@@
where the scaling $\hat{n}/N$ accounts for the fact that input photons are distributed across the $N$ detection bins and the second-order term proportional to $\gamma$ probes the presence of nonlinear absorption properties [58, 30]. Note that we probed different orders of the Taylor approximation in the variable $\hat{n}/N$ and found that a linear response function suffices for our data. However, as a concrete example, we here focus on the second-order nonlinear case. Together with Eq. (8), we have the function
@@
Figure 5: Determination of the response function [Eq. (10); blue, dashed lines] from measured mean click numbers [Eq. (8); thick, purple points] for $N=4$ and $8$ time-bin detection setups in the top and bottom depiction, respectively. In both cases, the response is governed by a linear behavior, with quantum efficiencies $\eta|_{N=4}=60.8\%$ and $\eta|_{N=8}=60.5\%$. Dark counts (less than $2\times 10^{-4}$) and nonlinear contributions ($\gamma\bar{n}^{2}/N^{2}<10^{-3}$) are negligible in both cases. For $N=4$ and $8$, the coefficient of determination is $R^{2}>0.9999$, quantitatively showing a near-perfect agreement between data and fit.
We use a nonlinear regression in Fig. 5 to retrieve the coefficients in Eq. (10). For both detector variants, $N=4$ and $8$, the dark count contributions are negligible, $\nu|_{N=4}=1.8\times 10^{-4}(1\pm 0.4)$ and $\nu|_{N=8}=5.8\times 10^{-6}(1\pm 8)$. Also, the nonlinear contribution that scales with $(\bar{n}/N)^{2}$ is small, $\gamma|_{N=4}=1.9\times 10^{-2}(1\pm 0.6)$ and $\gamma|_{N=8}=6.2\times 10^{-2}(1\pm 0.5)$. (None of these parameters is different from zero with statistical significance, a three-standard-deviation error margin.) Most importantly, the quantum efficiency of the click-counting device is about $60\%$, $\eta|_{N=4}=60.8\%(1\pm 0.004)$ and $\eta|_{N=8}=60.5\%(1\pm 0.005)$, including both optical losses in the multiplexing stage and non-unit efficiencies of individual on-off detectors.